Leg 205 was originally scheduled to begin in San Diego, California, on 6 September 2002. Because of the threat of a dock workers strike in the western United States, the port call was changed to Victoria, Canada. As a result, the start of the Victoria port call changed to 2 September.
Leg 205 officially began with the first line ashore at 0900 hr 2 September. A number of changes resulted from the short time between the decision to change the port call and the start of the port call: (1) the ODP technical staff crossover occurred on 4 September; (2) the Transocean staff crossover took place on 6 September; and (3) the Leg 205 science party boarded the ship on 17 September in Acapulco, Mexico, during the transit from Victoria, Canada, to proposed Site 1039R-A.
Prior to loading the Leg 205 supplies, a substantial amount of equipment used to support the Leg 204 gas hydrate objectives had to be taken off the ship. Pressurized gas hydrate samples, gas hydrate samples immersed in liquid nitrogen, and 340 boxes of core samples were off-loaded. A 20-ft refrigerated van used for processing and storing the gas hydrate samples on board was also removed from the ship. A number of special coring tools (HYACINTH [Deployment of HYACE Tools in New Tests on Hydrates]) and special laboratory equipment for gas hydrate research also were off-loaded.
Immediately upon arrival in Victoria, 1095 metric tons of marine gas oil was taken on board and we started to load 60 metric tons of barite and 80 metric tons of sepiolite. After the food for Leg 205 was loaded, the following equipment was loaded: (1) seven joints of 16-in casing, (2) 30 joints of 103/4-in casing, (3) 139 joints of 41/2-in casing, (4) a new spool of core winch wire, (5) four standard reentry cones, (6) four CORK-II wellheads and associated instruments, (7) four 16-in casing hangers, (8) two 103/4-in casing hangers, (9) a new bicentered hole opener, and (10) other miscellaneous hardware.
An essential element of the long-term geochemical observatories to be installed during the leg was the downhole water, gas, pressure, and temperature sampling/monitoring equipment. During the port call, six downhole OsmoSamplers were assembled and stored in a water bath in the downhole measurements laboratory so that they would be completely prepared for deployment later in the leg. During assembly, temperature sensors were placed inside of them. The pressure meters and data logger units were checked out and programmed in preparation for deployment.
At the request of the port authority, starting on 4 September, loading operations had to be suspended at 1700 hr each day for noise abatement. The ship left Victoria at 0754 hr on 7 September and began the transit to Acapulco, Mexico, to pick up the Leg 205 science party.
Because of Leg 204 operations continuing up to the last possible minute and a very short (39 hr) transit to Victoria, Leg 204 personnel had only minimal time for the routine end-of-leg maintenance of the drilling equipment; much of this activity took place during the transit. In addition, the new core winch wireline was installed and some minor plumbing runs associated with the CORK-II wellheads were made during the transit.
Calm seas were encountered during the initial portion of the transit. On 15 September, however, the vessel began passing through the edge of a tropical depression (36- to 42-nmi/hr winds and 20-ft seas) which reduced the ship's speed. The weather moderated during the day of 16 September, and moderate seas were encountered for the rest of the transit to Acapulco, Mexico.
The transit ended with the first line ashore at 1030 hr on 17 September in Acapulco, Mexico. The total transit was 2636 nmi and took 239.4 hr at an average speed of 11.0 nmi/hr.
The primary objective of the brief Acapulco port call was to board the Leg 205 scientists and load a few pieces of equipment that were essential for Leg 205 operations but did not make the Victoria port call. This included two 103/4-in casing hangers and the 32-in vibration isolated television (VIT) camera guide sleeve.
Because the weather forecast for the transit to Site 1039R-A did not look favorable for using the cranes while under way, we decided to assemble the Site 1039R-A reentry cone while in port. Work on the reentry cone began at 1200 hr. By 1700 hr the reentry cone had been bolted together and tack welded enough to make it secure for the transit, and a pilot was requested for departure. During this time, a required safety briefing on the vessel's survival gear was given to the science party. In addition, we also took advantage of the fact that the vessel was moored with starboard side to the dock and lowered lifeboats 2 and 4 to the water and started their engines. With all tasks completed in Acapulco, the last line was released at 1738 hr on 17 September and we began the transit to proposed Site 1039R-A.
The final welding of the reentry cone was completed (0600 hr on 18 September) during the transit to proposed Site 1039R-A. We also held additional safety meetings, introduced key Transocean personnel and the ODP operations manger to the science party, and gave science party members tours of the ship to familiarize them with the vessel.
At 1053 hr on 19 September, we changed course toward shore to disembark one of the ship's steward because of a family emergency. At 1830 hr on 19 September, the vessel was positioned just off the south breakwater at Puerto Madero, Mexico. A small fishing boat arrived from Puerto Madero and took the steward to shore. At 1845 hr the vessel resumed the transit to proposed Site 1039R-A.
At 1400 hr on 21 September, the vessel arrived at proposed Site 1039R-A.
The thrusters were lowered and the vessel was placed in dynamic positioning mode at 1431 hr on September 21, and we deployed a seafloor positioning beacon at 1620 hr. The upper guide horn was then laid out and the reentry cone moved to the center of the moonpool doors. Drill collars were picked up for the bottom-hole assembly (BHA), assembled, and stored in the derrick. Three joints of 16-in casing (42.5 m) were assembled, and a 16-in casing hanger was made up to the casing string. The casing running tool was then attached to the casing string, and the casing string was lowered into the reentry cone until the casing hanger latched into the reentry cone. The weight of the reentry cone was picked up to verify it had latched properly. We then released the casing running tool and pulled out of the reentry cone. A visual inspection confirmed that the casing hanger latch ring was properly seated in the reentry cone landing ring latch ring groove.
The next step was to attach a 143/4-in tricone bit to the stinger below the casing running tool. The rest of the BHA was then made up and lowered into the reentry cone until the running tool landed. The running tool was latched into the 16-in casing hanger, and the entire assembly (16-in casing, reentry cone, BHA) was picked up off the moonpool doors. The moonpool doors were opened, and at 2100 hr on 21 September, the reentry cone was lowered below the keel. The moonpool doors were then closed and the pipe trip to the seafloor began.
At 0745 hr on 22 September, with the 16-in casing shoe and bit near the seafloor, the jetting-in process began, initiating Hole 1253A. At 0945 hr, the reentry cone mud skirt landed on the seafloor and the running tool was released. This was verified using the VIT camera system. The seafloor depth was determined to be 4387.1 meters below rig floor (mbrf) (4376.3 mbsl), and the bottom of the 16-in casing shoe was at 43.82 mbsf. We then retrieved the drill string in preparation to assemble an RCB BHA.
After installing the reentry cone and 16-in casing, we assembled an RCB BHA with a 97/8-in bit and bit release. The bit was lowered to the seafloor, and Hole 1253A was reentered at 0555 hr 23 September. We then drilled ahead without coring from 4431.0 to 4444.0 mbrf (43.8 to 56.8 mbsf) where a DVTPP measurement was taken. The temperature and pressure data appeared to indicate that the tool moved while in the formation.
When the DVTPP was recovered, the colletted delivery system (CDS) was jammed with rust and sediment. This created a seal inside the upper barrel above the collets, trapping pressure inside the barrel. A vent hole was drilled in the upper barrel to prevent this from recurring, and the CDS was cleaned up and redressed with a new set of collets. After stroking the tool to verify performance, the CDS was set aside for use at a later time.
We continued to drill ahead without coring from 4444.0 to 4537.2 mbrf (56.8 to 150 mbsf) when the DVTPP was deployed a second time. Unfortunately, this run of the DVTPP also experienced tool movement while in the formation resulting in poor data. A review of the tool and deployment procedures is under way.
We then drilled ahead without coring to 4757.1 mbrf (369.9 mbsf), where we started RCB coring. Cores 205-1253A-1R through 4R, consisting of nannofossil chalk and claystone, were taken from 4757.1 to 4787.6 mbrf (369.9 to 400.5 mbsf). The rate of penetration (ROP) in the sediments was 23 m/hr, and we had a core recovery of 73%.
Cores 205-1253A-5R through 9R were cut from 4787.6 to 4817.0 mbrf (400.5 to 429.9 mbsf) in a very dense and hard gabbro sill. The ROP in the gabbro was 1.7 m/hr, and core recovery was 70.3%. Another change in the drilling occurred when we encountered sediments in Cores 205-1253A-10R through 13R. The ROP was 18.2 m/hr, and core recovery dropped to 17%.
We encountered more gabbro in Cores 205-1253A-14R through 28R from 4817.0 to 4916.0 mbrf (429.9 to 528.9 mbsf). A few thin sediment layers were present in the gabbro, but these were only minimally recovered. The ROP dropped to 1.4 m/hr, and core recovery rate was 50%.
At the end of cutting Core 205-1253A-25R, a small amount of erratic torque in the drill string was observed. With 53.2 total rotating hours on the bit, 45.2 hr of which were in hard rock, the decision was made to trip the drill string for a bit change. The bit cleared the seafloor at 1950 hr on 27 September. During the trip out of the hole, tight spots were noted at 4816, 4734, and 4647 mbrf. The bit cleared the rig floor at 0245 hr on 28 September. The used drill bit was in good condition but was slightly reduced in gage.
A new 97/8-in RCB bit was made up to the BHA. Because of the length of time drilling in hard rock, the mechanical bit released was also replaced. The BHA was spaced out for the RCB and lowered to the seafloor.
Hole 1253A was reentered at 1108 hr on 28 September. The bit was lowered into the hole to 4816 mbrf (429 mbsf), where excessive drag was encountered. We picked up the top drive and washed/reamed from there to the bottom of the hole at 4903.1 mbrf (516 mbsf). Once the 32 m of fill in the bottom of the hole was cleaned out, we resumed RCB coring.
After the bit change, we continued RCB coring from 4903.1 to a total depth of 4987.1 mbrf (515.9 to 600.0 mbsf) and recovered 59.6 m (71%). The penetration rate through this hard gabbro section was a painstakingly slow 1 m/hr. The lowermost 6 m of the hole appeared to have numerous ledges and may be deviated. Because of the slow penetration rate and no clear indication that we were close to the bottom of the gabbro, we decided to stop coring.
Next, we circulated 30 barrels of sepiolite to clean out the hole and made a wiper trip in preparation for logging. While lowering the pipe back to bottom during the wiper trip, an obstruction was encountered at 4920 mbsf (533 mbsf). The top drive was picked back up and used to rotate and wash to the bottom of the hole. Several ledges were reamed on the way back to total depth. Torque in the bottom 6 m of the hole led us to speculate that this section may be deviated. As this section is so close to the bottom of the hole, there was no way to confirm this with a survey.
After the 2 m of fill in the hole was washed out, 30 barrels of sepiolite was circulated to ensure all the debris was washed out of the hole. Then, in preparation for logging, the entire borehole volume hole was displaced with 185 barrels of sepiolite. The bit was pulled out of the hole, and the was ship offset 50 m north to drop the bit on the seafloor. At 2308 hr on 2 October, the pipe reentered Hole 1253A and the end of the pipe was positioned at the 16-in casing shoe at 4431 mbrf (44 mbsf).
The logging equipment was rigged up, and the triple combo tool string was lowered into the hole. The tools would not penetrate past a bridge at 4537 mbrf (150 mbsf). Because the upper sediment section had been logged previously during Leg 170 and our primary logging objective was the deeper part of the section, we decided to postpone logging until after we set the 103/4-in casing into the top of the first sill.
We retrieved the drill string, and the bottom of the BHA was back on the rig floor at 1615 hr on 3 October. To prepare for the installation of the 103/4-in casing, we needed to widen the 97/8-in hole to 143/4 in. We attached a 143/4-in reaming bit to the BHA, lowered it to the seafloor, and reentered Hole 1253A at 0044 hr on 4 October. We then drilled (without coring) from the 16-in casing shoe (4431 mbrf; 44 mbsf) to 23 m into the top of the upper gabbro (4810 mbrf; 422.8 mbsf). To make sure the hole was in the best condition possible to install the 103/4-in casing, the entire borehole volume was circulated two times, a wiper trip was made (without any significant hole problems), the borehole volume was circulated again, and then the hole was filled with sepiolite. The 143/4-in reaming bit was then pulled out of the hole; it reached the rig floor at 1345 hr on 5 October.
After assembling 30 joints of 103/4-in casing (411 m) and a cementing shoe, we lowered it to the seafloor and reentered Hole 1253A at 0210 hr on 6 October. We picked up the top drive with a cementing manifold and landed the 103/4-in casing hanger in the reentry cone ar 0445 hr on 6 October. The casing was cemented in place with 40 barrels of cement (15.8 lb/gal). The cementing was completed at 0650 hr, the casing running tool was released, the cementing manifold was laid out, the top drive was racked back, and the drill string was pulled out of the hole. The camera system was recovered while the pipe was pulled out of the hole, and the drill string was back on the rig floor at 1530 hr on 6 October.
Our next step was to drill out the cement plug at the bottom of the casing as well as clean out the hole to total depth (600 mbsf) in preparation for logging and installing the CORK. We assembled the BHA with a tricone bit (Reed HP21 97/8-in mill tooth), ran it down to the seafloor, and reentered the hole at 0011 hr on 7 October. The top of the cement was encountered at 4795 mbrf (407.8 mbsf), which is ~5 m above the bottom of the 10
The cementing plug and cement inside the casing were drilled out in 45 min, and then the hole was cleaned to 4987.1 mbrf (600.0 mbsf). The hole was swept clean with 30 barrels of sepiolite, and we started a wiper trip by raising the bit up to the 103/4-in shoe. On the way back down to 600 mbsf, 20,000 lb of weight was taken at 4976 mbrf (588.8 mbsf), so we rotated and washed the hole from there to total depth. We found 7 m of fill in the bottom of the hole. The hole was swept again with 50 barrels of sepiolite and then displaced with 60 barrels of sepiolite in preparation for logging.
Before we could log, we had to drop the bit on the seafloor. The bit was pulled clear of the reentry cone at 1630 hr on 7 October, and the vessel was offset 50 m north. The rotary shifting tool (RST) was deployed on wireline to release the bit. The mechanical bit release (MBR) shifted smoothly; however, the bit did not drop off. The RST was picked up and then dropped on top of the MBR support bearing, knocking the bit free at 1840 hr.
We reentered Hole 1253 (at 2027 hr), positioned the bottom of the pipe near the 103/4-in casing shoe at 4798 mbrf (410.8 mbsf), racked back the top drive, and started to rig up for logging. After assembling the triple combo tool string and lowering it into the hole to 4917 mbrf (530 mbsf), we lost communications with the tool and it had to be pulled out of the hole. The Hostile Environment Litho-Density Sonde was leaking oil; it was replaced with the Hostile Environment Litho-Density Tool. The reconfigured triple combo was lowered down the drill string until it landed on a bridge at 4917 mbrf (530 mbsf). The tools could not be worked past the bridge, so the hole was logged from that point up to the seafloor and then pulled out of the hole.
The FMS-sonic velocity tool string was assembled and lowered into the hole until it landed on a bridge at 4917 mbrf (530 mbsf). The tools would not pass the bridge, and we logged from there up to the 103/4-in casing shoe (4798 mbrf; 411 mbsf). The FMS was lowered for a second pass, and this time it passed through the bridge at 4917 mbrf, coming to rest on another bridge at 4951 mbrf (564 mbsf). We logged from that point up to the 103/4-in casing shoe and then pulled the tools out of the hole. Once we had disassembled the logging tools, we pulled the drill string out of the hole, clearing the reentry cone at 2215 hr on 8 October. After the bottom of the pipe was back on the rig floor (0515 hr on 9 October), the drill line was slipped and cut.
The next operational step at Hole 1253A was to deploy the CORK-II. After preparing the rig floor and all the tools required, the 41/2-in casing screen was picked up and hung off at the rotary table. A latch nipple for landing/latching the OsmoSampler inside the casing was attached to the screen. To the top of the screen, we attached one full joint of 41/2-in casing, one 2-m-long piece of 41/2-in casing, and one 3-m-long piece of 41/2-in casing. The screen was lowered into the moonpool, where the flat umbilical was connected to the two 1/4-in pressure monitoring hydraulic lines extending from the top of the screen. The screen was then lowered through the moonpool as the umbilical was strapped to the 41/2-in casing.
A CORK-II inflatable packer was picked up next and attached to the top of the 41/2-in casing. The packer was lowered into the moonpool, where the umbilical was severed and connected to the two 1/4-in pressure monitoring hydraulic pass through lines extending from the bottom of the packer. The packer was then lowered through the moonpool, and the umbilical was connected to the two 1/4-in pressure monitoring hydraulic pass through lines extending from the top of the packer.
We then attached 33 joints of 41/2-in casing, one 2-m-long piece of 41/2-in casing, and one 3-m-long piece of 41/2-in casing (467.66 m total length) to the top of the packer. The umbilical was strapped to the 41/2-in casing and centralizers attached to the casing as the screen/packer subassembly was lowered through the moonpool.
During deployment of the 41/2-in casing string and umbilical, at ~16 m above the packer, one of the three umbilical 1/4-in hydraulic lines was severed and split out of the umbilical. A 1/4-in pressure monitoring line termination screen was connected to the split-out umbilical hydraulic line. The termination screen will be used to monitor the pressure above the packer and below the 103/4-in casing shoe. The termination screen was banded to the 41/2-in casing, and centralizers were positioned at the top and bottom of the termination screen to protect it.
The CORK-II running tool was latched onto the CORK-II wellhead. The wellhead was then picked up and made up to the top of the 41/2-in casing string. The wellhead was lowered into the moonpool where the umbilical was severed and connected to the bottom of the wellhead. The packer setting hose was made up between the CORK-II running tool and the top of the wellhead, completing the assembly.
All of the sampling valves and bleed valves on the wellhead were opened, and the wellhead was lowered into the water to purge the hydraulic lines of air. The 1/16-in hydraulic lines connecting the pressure meter control valves to the pressure meters were filled with water prior before the wellhead was picked up. The wellhead was raised back to the moonpool level, where all valves were closed. Large rubber bands were then attached to the individual valve handles such that they would hold the valves in the closed position during the deployment. This was done to prevent the valves from partially opening during the deployment as happened with the ACORK valves deployed during Leg 196. With all hydraulic lines purged, all valves closed, and a last-minute inspection of the wellhead completed, the CORK-II assembly was lowered to the seafloor.
Hole 1253A was reentered at 0320 hr on 10 October. The 41/2-in casing screen shoe was lowered to 4874 mbrf (487 mbsf), where we picked up the top drive and lowered the wireline sinker bar assembly into the hole for a sounding run. The wireline sinker bar assembly reached 4915 mbrf (527.8 mbsf), which was 7 m below the OsmoSampler sinker bar's final position.
The lower weak link (900 lb) was made up to the top of the OsmoSampler sinker bar. This subassembly was then picked up and landed in the top of the drill string. The upper OsmoSampler was attached to the latch/running tool subassembly with 4.88 m of 3/8-in braided nylon rope. The upper weak link (1500 lb) was attached to the bottom of the upper OsmoSampler followed by 13.4 m of 3/8-in braided nylon rope. The lower OsmoSampler was then attached to the rope.
The Hole 1253A OsmoSampler subassembly (23.61 m over all length) was picked up, and the bottom of the lower OsmoSampler was connected to the lower weak link on top of the sinker bar that hung off in the drill pipe. The entire assembly (27.31 m overall length) was then lowered down the drill string until the latch/running tool subassembly landed on the drill string. The wireline sinker bar assembly was made up to the latch/running tool subassembly, and then the entire assembly was slowly lowered down the drill string in steps to prevent differential pressure from damaging the osmotic pumps.
The OsmoSampler assembly was lowered until the latch/running tool subassembly landed in the latch nipple on top of the 41/2-in casing screen. The latch/running tool was jarred down on to set and lock the latch (lock mandrel) in the latch nipple. A 3,000 lb overpull with the wireline confirmed that the latch was locked in place. The latch/running tool was then jarred up on, shearing the running tool release pin and releasing the running tool from the latch at 1110 hr on 10 October. The running tool and wireline were then pulled out of the hole.
The packer setting go-devil was then dropped down the inside of the drill string. The CORK-II wellhead was then lowered into the hole until it landed and latched into the reentry cone. A 5,000 lb overpull with the drill string confirmed that the wellhead was latched into the reentry cone.
The drill string was pressured up to 800 psi and held for 30 min to set the packer. While the drill string was being pressurized, the pressure gauge fluctuated once as if there was a sudden small change in volume. Also, occasionally while holding the 800-psi pressure, the pump had to be engaged at 23 strokes per minute to maintain the pressure. After 30 min, the drill string pressure was increased to 1,800 psi and held for 10 min to activate the spool valves (to connect the pressure sensors in the wellhead to the downhole screens). Finally, all of the drill string pressure was bled off through the rig floor standpipe manifold relief valve.
The final step was to install the ROV platform around the wellhead on top of the reentry cone. The landing platform and deployment bridle were assembled in the moonpool and then lowered down the drill string to the reentry cone on the logging line. After landing the platform on top of the reentry cone the deployment bridle acoustic releases were activated at 2245 hr on 10 October. The logging line was raised and the weight indicator showed a reduction in weight of 800 lb, confirming that the platform had been released. The logging line and platform deployment bridle were then pulled out of the hole.
We then deployed the camera system to visually inspect the installation. No problems were observed during the visual inspection, so the CORK-II running tool was released from the wellhead at 0050 hr on 11 October.
The wireline sinker bar assembly was deployed to recover the packer setting go-devil to prevent having to retrieve the drill string when it was full of water. However, the go-devil was encountered at 400 mbrf instead of at the CORK-II running tool, where it should have been. It took 1.5 hr and jarring with 8,000 lb of overpull to recover the go-devil. When it was recovered, the go-devil seals were missing. The top drive was racked back, and the drill string and camera system were pulled out of the hole. The CORK-II running tool cleared the rig floor at 0945 hr on 11 October.
While the drill string was being retrieved, the ship moved in dynamic positioning mode to Site 1254 (1.1 nmi). The beacon from Site 1253 was recovered at 0430 hr on 11 October and redeployed at 0945 hr on 11 October at Site 1254A.
After the drill string was pulled out of Hole 1253A, the ship moved in dynamic positioning mode to Site 1254 (1.1 nmi) while the drill string was being retrieved. The beacon from Site 1253 was recovered at 0430 hr on 11 October and redeployed at 0945 hr on 11 October at Site 1254A.
The first operation at Site 1254 was to install a reentry cone and 16-in surface casing. Two joints (29.37 m) of 16-in casing were made up to a casing hanger and latched into a standard reentry cone. The reentry cone was lowered to the seafloor and Hole 1254A was initiated at 2225 hr on 11 October. The water depth was determined to be 4194 mbrf.
The sediments proved to be quite consolidated near the seafloor, and 10 hr was required to jet in the 16-in casing. The casing running tool was released at 0850 hr on 12 October, and the drill string was pulled out of the hole clearing the reentry cone at 0900 hr. The camera system was recovered, and the casing running tool cleared the rotary table at 1645 hr 12 October.
After installing the reentry cone and 16-in casing, we assembled a RCB BHA and lowered it to the seafloor. Hole 1254A was reentered at 0110 hr on 13 October, and we drilled without coring from the 16-in casing shoe at 4223 mbrf (29 mbsf) to 4244 mbrf (50 mbsf). A DVTPP measurement was made at 50 mbsf, but the data were not usable. Hole 1254A was then drilled from 4244 to 4344 mbrf (50 to 150 mbsf), where a second DVTP deployment was made; the data were once again not usable. Hole conditions through this part of Hole 1254A were stable, and the ROP was 9.3 m/hr.
RCB Cores 205-1254A-1R through 6R were cut from 4344 to 4397.8 mbrf (150 to 203.8 mbsf). A third DVTPP measurement was made at 203.8 mbsf, and for the first time this leg, good temperature data were recorded. The pressure data, however, were not good. RCB Cores 205-1254A-7R and 8R were taken from 4397.8 to 4417 mbrf (203.8 to 223.0 mbsf). Cores 205-1224A-1R through 8R penetrated 73 m, and we recovered 66.59 m (91%); the overall penetration rate was 10.8 m/hr. No hole problems were encountered, and the hole remained stable while coring this interval.
Hole 1254A was then drilled without coring from 4417 to 4494 mbrf (223 to 300 mbsf) in 8.5 hr (9 m/hr). The hole remained stable through this drilled interval requiring only periodic circulation of sepiolite mud to keep the hole clean.
To determine the upper and lower boundaries of the décollement, Hole 1254A was cored from 4494 to 4561.7 mbrf (300 to 367.5 mbsf). Cores 205-1254A-9R through 16R recovered 58.41 m (87%), and the penetration rate through this interval was 10 m/hr.
Sepiolite mud was circulated periodically over this cored interval to maintain hole stability; however, none of the anticipated hole problems associated with drilling through a décollement occurred. The hole was displaced with 111 barrels of sepiolite, and the RCB bit was pulled out of the hole. The RCB bit cleared the reentry cone at 0530 hr on 16 October and reached the rig floor at 1230 hr.
A BHA was assembled with a 143/4-in bit for opening the hole in preparation for installing the 103/4-in casing string. We reentered Hole 1254A at 2135 hr on 16 October and rotated and washed from the 16-in casing shoe at 4223 mbrf (29 mbsf) to 4527 mbrf (333 mbsf). Fifty barrels of sepiolite mud were then circulated to clean out the hole, and then we circulated twice the borehole volume. We then conducted a wiper trip that encountered a tight spot at 4395 mbrf (201 mbsf). The bit was washed and reamed through the tight spot and lowered back down to the total depth of the 143/4-in hole at 4527 mbrf (333 mbsf); 2 m of fill at the bottom of the 143/4-in section of the hole had to be washed out.
We circulated another 50 barrels of sepiolite and twice the hole volume with seawater. A second wiper trip was undertaken back to the 16-in casing shoe and encountered tight spots at 4482, 4420, and 4386 mbrf (288, 226, and 192 mbsf). During the wiper trip back down, the bit encountered a bridge at 4386 mbrf (192 mbsf); it could support ~20,000 lb of drill string weight. We washed and reamed through this tight spot and back down to 4527 mbrf (333 mbsf).
Once again, we swept the hole with 60 barrels of sepiolite and twice the hole volume with seawater. The hole was then displaced with 255 barrels of sepiolite to help stabilize the hole while we prepared to assemble and install the 103/4-in casing string. The drill string encountered no significant drag while pulling out of the hole. The bit cleared the reentry cone at 2215 hr on 17 October, and the bit reached the rig floor at 0500 hr on 18 October.
Twenty-eight joints (324 m) of 103/4-in casing were made up and attached to a 103/4-in casing hanger. The cementing tools were assembled and attached to the casing running tool which in turn was made up to the 103/4-in casing string. The 103/4-in casing string was then lowered to the seafloor, and Hole 1254A was reentered at 1738 hr on 18 October.
The 103/4-in casing shoe was lowered to 4426 mbrf (232 mbsf) when it suddenly appeared to hang up as identified by reduction in string weight of 20,000 lb. The drill string was immediately picked up to the hanging weight of the string. The casing was lowered again and took weight at the same depth.
A stand of drill pipe had just been picked up, and there was not much traveling space in the rig available to work the casing up and down. The casing was worked within these traveling limits without being able to pass the tight spot.
The space-out was such that the drill string could not be picked up without activating the Crown-O-Matic (the stand of drill pipe was too close to the top of the derrick) and could not be lowered and secured on the rig floor without putting the casing in compression. Thus the ship was offset far enough to allow the drill string to be landed on the drilling elevator stool without putting the casing in compression. A stand of drill pipe was laid out, and the ship was moved back over the hole. When we raised the drill string so that we could remove a single joint of drill and attach the top drive, the reentry cone was pulled up off the seafloor along with the 103/4-in casing string.
The 103/4-in casing string and reentry cone were worked in an unsuccessful attempt to free the reentry cone from the casing string. The VIT camera system was lowered to the reentry cone for a visual inspection. It appeared that a joint of 103/4-in casing had collapsed inside the reentry cone. With all options for freeing the reentry cone from the 103/4-in casing string exhausted, the 103/4-in casing string, with the reentry cone and 16-in casing attached, was pulled out of the hole.
Once the casing string with the stuck reentry cone was clear of the seafloor, we offset the ship 0.5 nmi perpendicular to the seismic line on which Site 1254 is located. If the reentry cone or other hardware fell off during recovery, we didn't want it to fall in the area where the scientists wanted to attempt another CORK-II installation. The reentry cone was pulled into the moonpool at 0845 hr on 19 October. While laying out the 103/4-in casing string, we found four joints that were bent. This bending likely occurred during attempts to break off the casing inside the reentry cone at the seafloor. Approximately 3 m of a collapsed joint of 103/4-in casing was found protruding out of the top of the reentry cone.
It was deemed unsafe, and too costly in time, to attempt to salvage the reentry cone and casing. Thus preparations were made for jettisoning the reentry cone. The connection between joints 7 and 8 was partially loosened. The reentry cone was then lowered below the keel of the ship and hung off near the top of joint 9 in slips at the rotary table. Although the casing was hung off in the slips, the elevator was left attached to the casing and positioned just beneath the collar. A padeye was then welded to the casing just below the elevator to prevent the joint from jumping upward when the reentry cone was released. With all personnel cleared from the rig floor and the moonpool area, the rotary table was engaged to finish unscrewing the loosened connection, and at 0920 hr on 19 October the reentry cone was jettisoned, ending Hole 1254A.
The scientists decided to make a second attempt at installing a CORK-II at Site 1254. The ship was then moved to ~48 m at N36°E from Hole 1254A. After the 103/4-in casing handling equipment was rigged down, the 16-in casing handling equipment was rigged up. A reentry cone, which had been previously assembled for our next site, was moved into the well center on top of the moonpool doors. Two joints of 16-in casing (20 m) plus a 16-in casing cementing shoe were then made up to a 16-in casing hanger, and the assembly was latched into the reentry cone. The jetting BHA was then made up and latched into the 16-in casing hanger. The reentry cone was then picked up off the moonpool doors, and the reentry cone was then lowered to the seafloor. We started washing in the 16-in casing at 2550 hr 19 October, beginning with Hole 1254B. After it was washed into the seafloor to 20.42 mbsf in 6.5 hr, the casing running tool was released and the drill string was retrieved. Our next step was to drill a 143/4-in hole into which we could install the 103/4-in casing. We assembled a drilling BHA with a 143/4-in bit, lowered it to the seafloor, and reentered Hole 1254B at 2134 hr 20 October.
The 143/4-in hole was advanced to 4418 mbrf (231 mbsf) when the upper fault zone was penetrated. As we advanced through the fault, increased fill and torque were encountered. The fault had to be reamed and swept clean by circulating sepiolite before the bit could penetrate deeper.
When the bit reached 4452 mbrf (265 mbsf), the hole began to pack off around the drill string. The bit was raised 20 m and then washed and reamed back down to 265 mbsf. We drilled to 4465 mbrf (278 mbsf) when the pipe became stuck. After three hours of working the pipe with overpull up to 200,000 lb, the pipe could be raised again.
Because of the poor drilling conditions, we decided to conduct a wiper trip. The bit was pulled up to 4289 mbrf (102 mbsf), encountering 80,000 lb drag from 4453 to 4415 mbrf (266 to 228 mbsf). It took 6.75 hr to wash and ream back to the bottom of the hole at 4465 mbrf (278 mbsf); 40 m of fill in the bottom of the hole had to be drilled/washed out. With hole conditions deteriorating and no hope of setting casing to the décollement, we decided to install a CORK-II to monitor the upper fault zone.
The bit was pulled up to the 16-in casing shoe at 4210 mbrf (23 mbsf), and then the pipe was lowered back down. The bit encountered resistance (took weight) at 4397 mbrf (210 mbsf) and was washed and reamed from there to 4424 mbrf (237 mbsf). With the bit at the bottom of the hole, we circulated bottoms up in preparation for setting 103/4-in casing. Because the 143/4-in hole extended through the zone of interest, the hole was left full of seawater and no sepiolite mud was used. The bit was pulled out of the hole and was back on the rig floor at 0830 hr on 23 October.
Seventeen joints (197 m) of 103/4-in casing were then assembled, and a casing hanger was attached to the top. The cementing and casing running tools were latched into the casing hanger, and the entire assembly was lowered to the seafloor. At 1822 hr on 23 October, Hole 1254B was reentered with the 103/4-in casing string. It appeared that the hole was producing fluid as we observed, as evidenced by a billowing cloud within the reentry cone and water flowing out of the drill pipe at the rig floor between connections.
The casing was lowered to 4360 mbrf (173 mbsf) when it encountered some resistance and began to take weight. At this point, we picked up the top drive to wash the casing in. The cementing manifold was picked up, and the casing was landed at 2045 hr on 23 October. We confirmed that it was latched in and cemented it in place with 20 barrels of 15.8 lb/gal cement. The 103/4-in casing shoe was at 198.99 mbsf. The casing running tool was released and was back on the rig floor at 0915 hr on 24 October.
Our next operation was to drill out the cement plug and clean out a rathole for the CORK. We assembled a drilling BHA with a 97/8-in tricone bit, lowered it to the seafloor, and reentered Hole 1254B at 1716 hr on 24 October. After lowering the bit to the casing shoe, it took 1.5 hr to drill through the shoe. The hole was then cleaned to 4424 mbrf (237 mbsf) in preparation for deploying the CORK-II. The hole was then circulated bottoms up and left with only seawater in it. The drill bit was pulled out of the hole and was back on the rig floor at 0445 hr on 25 October.
After some routine preventative maintenance (slipping and cutting the drill line), then we started assembling the CORK-II. A screen, packer, 15 joints of 41/2-in casing (230 m), umbilical, centralizers, and the wellhead made up the complete CORK-II. The wellhead was lowered to the moonpool doors, where the pressure data logger was tested; it was then lowered into the water to purge the hydraulic lines. At 1445 hr on 25 October, the CORK-II assembly was lowered to the seafloor and Hole 1254B was reentered with the CORK-II at 2050 hr. The 41/2-in casing was lowered to 4366 mbrf (179 mbsf) when it encountered some resistance and began to take weight. The top drive was picked up so that we could jet the 41/2-in casing in. The jetting process was hampered as a result of having to maintain the drill string pressure below 400 psi to prevent inflation of the packer.
After ~8 hr of attempting to jet the 41/2-in casing in with the active heave compensator on at 0745 hr on 26 October, the top joint pin of the 41/2-in casing failed through the last engaged thread. With the CORK-II wellhead still attached to the running tool, the vessel was offset to observe the reentry cone. At the time of the failure the end of the pipe was ~10 m above the reentry cone. The broken end of the 41/2-in casing was extending out of the reentry cone, but otherwise the reentry cone did not appear to have any observable problems. The CORK-II wellhead was retrieved and was back on the rig floor at 1645 hr on 26 October, ending Hole 1254B. The scientists decided to abandon efforts to deploy a borehole installation at this site and decided to move to Site 1255 (proposed Site 1043R).
The Site 1254 beacon was released at 1106 hr and was back on deck at 1221 hr on 26 October. The ship moved the ~1.1 nmi to Site 1255 in dynamic positioning mode while the CORK-II wellhead was being recovered.
The beacon at Site 1254 was released at 1106 hr and was on deck at 1221 hr on 26 October. The ship moved the ~1.1 nmi to Site 1255 in dynamic positioning mode while the CORK-II wellhead was being recovered. We deployed a seafloor position beacon at Site 1255 at 2116 hr on 26 October.
A reentry cone was then moved onto the moonpool doors and two joints (20 m) of 16-in casing were assembled to a 16-in casing hanger and latched into the reentry cone. A jetting BHA was assembled and latched into the 16-in casing hanger. At 2000 hr on 26 October, the reentry cone was lowered to the seafloor. Because of complex seafloor topography in the vicinity of the drill site, the 3.5-kHz precision depth recorder returned six different echoes. The most conservative shallowest depth indicated (4234 mbrf) was used. The casing was lowered to 4325 mbrf before the seafloor was detected by a reduction in drill string weight. Hole 1255A was started at 0550 hr on 27 October when jetting-in of the 16-in casing began.
The 16-in casing was jetted in to 4343.49 mbrf (20.49 mbsf) in 5 hr. The casing running tool was released, pulled out of the hole, and was back on the rig floor at 2045 hr on 27 October.
The next step was to prepare the hole for installing 103/4-in casing. A BHA with a 143/4-in drilling bit was assembled, lowered to the seafloor, and Hole 1255A was reentered at 0430 hr on 28 October. We drilled from 4343.49 to 4446 mbrf (20.49 to 123 mbsf). The hole was swept with 50 barrels of sepiolite, and then we circulated the complete borehole volume with seawater. We then conducted a wiper trip by raising the bit to the 16-in casing shoe at 4343.49 mbrf (20.49 mbsf) and then lowering it back down to 4429 mbrf (106 mbsf), where slight weight was taken. The hole was washed and reamed from 4429 mbrf (106 mbsf) back to total depth (4446 mbrf; 123 mbsf). The hole was again swept with 50 barrels of sepiolite, circulated twice the hole volume with seawater, and 85 barrels of sepiolite pumped into the hole in preparation for deploying 103/4-in casing. The bit was then pulled out of the hole and was back on the rig floor at 0526 hr on 30 October.
Ten joints of 103/4-in casing (115.5 m) were made up with 103/4-in Dril-Quip casing hanger. A casing cementing plug was then made up to the Dril-Quip casing running tool that, in turn, was latched into the casing hanger. The 103/4-in casing string was then lowered to the seafloor. Hole 1255A was reentered with the casing at 1629 hr on 29 October. The casing was lowered to 4425 mbrf (102 mbsf) when slight resistance was encountered. The top drive was picked up and the casing was washed in to a total depth of 117.42 mbsf. After it was confirmed that the casing hanger had latched into the reentry cone (10,000 lb overpull), the casing was cemented in place with 20 barrels of 15.8-lb/gal cement.
The casing running tool was released at 1940 hr on 29 October and pulled clear of the reentry cone. Before the drill string was tripped back to the ship, the drill line had to be slipped and cut. The casing running tool was back on deck at 0405 hr on 30 October.
An RCB BHA was assembled and lowered to the seafloor, and Hole 1255A was reentered at 1207 hr on 30 October. It took 4.5 hr to drill out the 103/4-in casing shoe with the RCB bit before coring began.
We then cut four RCB cores (Cores 205-1255A-1R through 4R) from 4446.0 to 4480.0 mbrf (132.7 to 157.0 mbsf) and recovered 7.22 m (21%). Once the base of the décollement was identified (based on a substantial change in penetration rate), we began preparing the hole for installing the CORK-II and OsmoSampler.
We swept the hole with 50 barrels of sepiolite, and the borehole volume was circulated with seawater. The bit was raised up to the 103/4-in casing shoe and then lowered back down to 4467 mbrf (144 mbsf), where the bit took weight; the hole was reamed and washed from there to 4476.0 mbrf (153.0 mbsf). Once again, we swept the hole clean with 50 barrels of sepiolite and then circulated the borehole volume with seawater. The bit was again raised up to the 103/4-in casing shoe and then lowered back down to total depth with only slight drag near the bottom of the hole. To keep the hole open without exposing the décollement to barite-weighted mud, 20 barrels of sepiolite was placed in the bottom of the hole. The bit was pulled up to 4416 mbrf (93 mbsf) inside the 103/4-in casing, and the hole was displaced with 32 barrels of heavy mud. The bit was pulled out of the hole and was back on the rig floor at 2015 hr on 31 October.
A CORK-II assembly was assembled consisting of a screen, packer, 123.4 m of 41/2-in casing, and a CORK-II wellhead. The CORK-II assembly was lowered to the seafloor to just above the reentry cone. The top drive was picked up so we could pump seawater through the end of the 41/2-in casing as it was lowered in the hole to displace the heavy mud and keep the screen clean. Hole 1255A was reentered at 1000 hr on 1 November and the 41/2-in casing was lowered to 4465 mbrf (142 mbsf) while seawater was circulated.
The OsmoSampler seat was attached to the coring wireline and lowered down the drill string. The seat landed and was jarred on to latch/lock the lock mandrel in place. When the wireline was recovered the seat was still attached and the running tool was partially sheared and the seat seals were missing. The running tool was redressed, new seals were installed on the seat, and the seat was lowered back down the drill string. The seat landed in what appeared to be the upper latch nipple at top of the packer. The mud pump was engaged, the drill string was pressured to 300 psi, and the seat passed through the upper latch nipple. The seat was then lowered until it landed in the middle latch nipple at top of the screen. Once again the pump was engaged, and the seat passed through the middle latch nipple. The seat was then lowered until it landed in the latch nipple inside the screen. The pump was engaged to fully latch the seat. The seat was jarred down on to latch the lock mandrel in place. An overpull of 3000 lb was applied and held, indicating that the lock mandrel was latched in place. The running tool was then jarred up on and released. The running tool was then pulled out of the hole and redressed.
When the OsmoSampler was attached to the running tool on the rig floor, the screws holding the OsmoSampler top connection in place were sheared. The running tool was removed, and the broken OsmoSampler was replaced with another one. The OsmoSampler fill port pipe plug (inside the top connection of the OsmoSampler) was protruding above the bottom of the blind box, preventing the double pin crossover sub to be properly attached. Bakerlok was applied to the double pin sub, and the sub was made up to the OsmoSampler top connection. Bakerlok was also applied to the collet housing, and the housing was made up to the OsmoSampler probe tip.
The OsmoSampler assembly picked up with a tugger and landed in the top of the drill pipe. The coring wireline was attached to the assembly, and it was lowered down the drill string in stages to allow pressure equalization of the osmotic pumps. Because the OsmoSampler lock mandrel had no seals, it easily passed through the upper latch nipple and was landed on top of the seat.
With the OsmoSampler probe tip stung through the seat, all of the OsmoSampler sampling ports were sealed off from the 41/2-in casing above. The pump was engaged to fully seat the OsmoSampler and help hold the assembly down during jarring. The running tool was jarred down on to latch the lock mandrel in place. When the wireline was raised, no overpull was observed and the weight of the OsmoSampler was lost, indicating that the running tool had released.
The OsmoSampler seat allows the pressure below and above the ported seal bore and plug to equalize. Thus, when the OsmoSampler probe tip is stung into the seat, the OsmoSampler is pressure balanced, so confirmation of latch-in (requiring pulling the OsmoSampler and deploying it again) was not critical to the installation. Once we verified that the OsmoSampler was at the correct depth (by tagging it with the wireline), we pulled the running tool out of the hole.
The modified packer setting go-devil was then dropped down the drill string and chased with the coring wireline without latch. The go-devil was tagged inside the CORK-II running tool, and the wireline was pulled out of the hole.
The CORK-II wellhead was then landed and latched into the reentry cone at 2000 hr on 1 November. The drill string was pressurized to 1,200 psi for 30 min to inflate the packer. The drill string pressure was then increased to 1,800 psi for 10 min to shift the spool valves. While the packer was set, the VIT camera system was pulled out of the hole in preparation for deploying the ROV platform.
The ROV platform was assembled around the drill string in the moonpool. The deployment bridle and the logging line were then attached. The platform was lowered down the drill string on the logging line and landed on the reentry cone.
At 0155 hr on 2 November, the acoustic releases were activated. When the logging line was picked up, overpull was observed, suggesting that the ROV platform was not completely free. The logging line was lowered and the acoustic releases were activated again. When the logging line was picked up, overpull was observed again. The overpull was increased above the weight of the platform, indicating the deployment bridle was hung up on something. The logging line was worked for 30 min without freeing the bridle when a straight pull was applied. The bridle came free with ~4500 lb overpull (9100 lb total wire load) and was pulled out of the hole.
When the bridle was recovered, one of the nylon strap termination eyes was found to have failed. The acoustic release yoke may have hung up in the nylon strap eye and had been pulled down to the platform strap bail. When overpull was applied, the platform was pulled up on one side only until the strap failed. If this scenario is correct, there should be no damage to the wellhead. After recovering the platform deployment bridle, the camera was lowered to inspect the installation and no problems with the installation were observed. The CORK-II running tool was released at 0615 hr on 2 November. The running tool and VIT were then recovered; the running tool was back on the rig floor at 1600 hr on 2 November.
Once the seafloor positioning beacon was released (0759 hr) and back on deck (0929 hr on 2 November), we moved back to Hole 1254B in dynamic positioning mode to attempt to fish the screen, packer, and 41/2-in casing that was left in the hole.
After finishing operations at Site 1255, we moved back to Hole 1254B in dynamic positioning mode to attempt to fish the screen, packer, and 41/2-in casing that was left in Hole 1254B.
The vessel was positioned over Hole 1254B at 1721 hr on 2 November. A fishing BHA was assembled, consisting of a choker loop formed from wire rope slings, a rig-fabricated jet sub, five 81/4-in drill collars, one tapered drill collar, and the usual two stands of 51/2-in drill pipe and a crossover sub.
The fishing tool was lowered to 4157 mbrf, and at 2343 hr 2 November the 41/2-in casing was snared. We applied 20,000 lb of overpull to the fish (41/2-in casing), and it suddenly dropped off. When the snare was set, the end of the wire rope slings was too far below the end of the BHA for us to visually confirm if the fish was still there. Also, the total weight of the casing string being fished was too small to be distinguished on the weight indicator. The vessel was offset 0.5 nmi as the BHA was pulled out of the hole so that if the fish dropped off, it would not land back on the reentry cone. When the fishing BHA was back on the rig floor, the snare was empty and the casing was not recovered. When the ship was positioned back over the hole so that we could inspect the reentry cone with the VIT camera, we observed that the casing was still sticking out of the reentry cone. After the camera was back on deck and all equipment secured we began the transit to Balboa, Panama, at 1600 hr on 3 November. Leg 205 ended with the last line ashore at 1030 hr on 6 November.
Next Section | Table of Contents